One criticism of racial profiling studies is that people’s driving frequency is often unaccounted for. This is a problem because, all else being equal, people who spend more time on the road are more likely to get pulled over eventually. The following table contains PPCS data narrowed down to black male respondents. The variables measure driving frequency and whether these respondents had been stopped by police for traffic offenses within the past 12 months. With an alpha of .01, conduct a five-step hypothesis test to determine if the variables are independent.

Driving Frequency Yes No Raw Marginal
Almost Every Day 214 946 1,160
Often 32 238 270
Rarely 6 360 366
Column Marginal 252 1,544 N=1,796
Answer #1

We have to perform Chi-square test for independence of two random variables (driving frequency and stop by police).

We have to test for null hypothesis \tiny H_0:\text{Driving frequency and stop by police are independent}

against the alternative hypothesis Hi : Driving frequency and stop by police are not independent

Observed frequencies \tiny \left(f_{ij}\right) are as follows.

Row total Observed frequencies Driving Almost everyday frequency Often Rarely Column total Stopped by police Yes No 214 946 3

Under null hypothesis, we obtain expected frequencies \tiny \left(e_{ij}\right) by multiplying corresponding row total and column total and dividing it by grand total as follows.

Stopped by police Expected frequencies Row total Yes No Almost everyday 162.76 997.24 1160.00 Driving Often 37.88 232.12 270.

Our Chi-square test statistic is given by

\tiny \chi^2=\sum_{i=1}^{m}\sum_{j=1}^{n}\frac{\left(f_{ij}-e_{ij}\right)^2}{e_{ij}}

Here,

Number of rows \tiny m=3

Number of columns \tiny n=2

Corresponding calculations (chi square component for each cell, which are to be added later) are as follows.

Row total Chi_square Stopped by police Yes No Almost everyday 16.13 2.63 Driving Often 0.91 0.15 frequency Rarely 40.06 6.54

\tiny \therefore \chi^2_{calculated}=\sum_{i=1}^{3}\sum_{j=1}^{2}\frac{\left(f_{ij}-e_{ij}\right)^2}{e_{ij}}=66.42

Degrees of freedom v = (m 1) (n − 1) = 2

\tiny \therefore \text{p-value}=P\left(\chi^2_{2}>66.42\right)=3.774758*10^{-15} [Using R-code '1-pchisq(66.42,2)']

Level of significance a = 0.01

We reject our null hypothesis if \tiny \text{p-value}<\alpha

Here, we observe that \tiny \text{p-value}=3.774758*10^{-15}<0.01=\alpha

So, we reject our null hypothesis.

Hence, based on the given data we can conclude that there is significant evidence that driving frequency and stop by police are not independent.